Preyer Corporation produces and sells a single product. Data concerning that product appear below:

Selling price per unit $ 250
Variable expense per unit $ 47.50
Fixed expense per month $ 506,493


The break-even in monthly dollar sales is closest to:

Noreen 4e Recheck 2017-16-03

$2,665,753

$492,407

$625,300

$1,153,501

Tribley Inc. has an operating leverage of 5.2. If the company's sales increase by 10%, its net operating income should increase by about:

40.4%

52.0%

1.9%

10.0%

Closser Corporation produces and sells two products. In the most recent month, Product M50S had sales of $36,000 and variable expenses of $11,280. Product H50G had sales of $49,000 and variable expenses of $17,620. The fixed expenses of the entire company were $46,110. The break-even point for the entire company is closest to: (Round your intermediate calculations to 2 decimal places and final answers to the nearest dollar amount.)

$75,010

$46,110

$69,864

$69,824

Data concerning Marchman Corporation's single product appear below:

Per Unit Percent of Sales
Selling price $ 160 100 %
Variable expenses 80 50 %
Contribution margin 80 50 %

The company is currently selling 6,800 units per month. Fixed expenses are $488,800 per month. Consider each of the following questions independently.

This question is to be considered independently of all other questions relating to Marchman Corporation. Refer to the original data when answering this question.

Management is considering using a new component that would increase the unit variable cost by $2. Since the new component would increase the features of the company's product, the marketing manager predicts that monthly sales would increase by 180 units. What should be the overall effect on the company's monthly net operating income of this change?

Decrease of $440

Increase of $14,400

Increase of $440

Decrease of $7,200

Darrow Company uses a predetermined overhead rate based on direct labor-hours to apply manufacturing overhead to jobs. Last year, the company worked 23,000 direct labor-hours and incurred $149,500 of actual manufacturing overhead cost. If overhead was underapplied by $8,740, the predetermined overhead rate for the company for the year must have been:

$6.12

$6.50

$6.88

$7.26

Sweet Corporation applies overhead to jobs on the basis of 70% of direct labor cost. If Job 107 shows $84,000 of manufacturing overhead applied, how much was the direct labor cost on the job?

$120,000

$58,800

$84,000

$109,200

Nagle Inc. uses a job-order costing system in which any underapplied or overapplied overhead is closed out to cost of goods sold at the end of the month. In October the company completed job O43G that consisted of 13,400 units of one of the company's standard products. No other jobs were in process during the month. The job cost sheet for job O43G shows that the total cost for the job was $720,250. During the month, the actual manufacturing overhead cost incurred was $176,460 and the manufacturing overhead cost applied to job O43G was $190,900. Also during the month, 7,400 completed units from job O43G were sold. No other products were sold. The cost of goods sold that would appear on the income statement for October is closest to:

$720,250

$383,310

$412,190

$705,810
Answer #1

Break even sales of Preyer Corporation

BES = Fixed cost/ Contribution margin ratio

Contribution margin ratio = contribution margin / sales = (250-47.5)/$250 =202.50/250 =81%

Break even sales = $506493/81% = $625300 (Third option is correct)